OpenStudy (gabbyalicorn):

The area of the base of a prism is 50 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism?

OpenStudy (gabbyalicorn):

50 x 20 x 7 x 2 ?

OpenStudy (michele_laino):

the area of both bases is: 50*2=...

OpenStudy (gabbyalicorn):

100

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now the area of lateral surface is: 30*7=...

OpenStudy (gabbyalicorn):

210

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so the total area is: 100+210=...

OpenStudy (gabbyalicorn):

310

OpenStudy (michele_laino):

that's right!

OpenStudy (gabbyalicorn):

:D

OpenStudy (michele_laino):

:)

OpenStudy (gabbyalicorn):

Yay! Thank you!

OpenStudy (michele_laino):

:)